OpenStudy (anonymous):

find dy/dx for y = 3cos(x) + csc(x)

OpenStudy (anonymous):

Do you know your derivative rules for trig functions? If not, you can find some here: http://www.sosmath.com/calculus/diff/der03/der03.html

OpenStudy (anonymous):

Heere is also a list of basic rules. This is a simple example where you have two functons added together. http://tutorial.math.lamar.edu/pdf/Common_Derivatives_Integrals.pdf

OpenStudy (anonymous):

okay so do i just ignore the fact that its dy/dx because the y is already isolated?

OpenStudy (anonymous):

Can you tell me what \[\Large \frac{ dy }{ dx } 3\cos(x)= ?\]

OpenStudy (anonymous):

-3sin(x)

OpenStudy (anonymous):

Ok and what is \[\Large \frac{dy}{dx} \csc(x)? \]

OpenStudy (anonymous):

-csc(x)cot(x)

OpenStudy (anonymous):

And when you get that, you add the two together. When you do a derivative of something in the dorm of \[\Large y=f(x)\] You actually have \[\Large \frac{ dy }{ dx }= f'(x)\] This is known as implicit differentiation. You will learn it later on. It is the underlying concept as to what you are doing.

OpenStudy (anonymous):

So your final answer is -3sin(x)-csc(x)cot(x)

OpenStudy (anonymous):

okay so then its just -3sin(x)-csc(x)cot(x)

OpenStudy (anonymous):

It's actually -3sin(x) + (-csc(x)cot(x)) but a positive plus a minus is a minus

OpenStudy (anonymous):

thank you very much !!

OpenStudy (anonymous):

Great job! Np